OpenStudy (experimentx):

Can anyone explain this to me? The fraction 49/98 is a curious fraction, as an inexperienced mathematician in attempting to simplify it may incorrectly believe that 49/98 = 4/8, which is correct, is obtained by cancelling the 9s. We shall consider fractions like, 30/50 = 3/5, to be trivial examples. There are exactly four non-trivial examples of this type of fraction, less than one in value, and containing two digits in the numerator and denominator. If the product of these four fractions is given in its lowest common terms, find the value of the denominator.

OpenStudy (ghazi):

49/98 = 4/8 could be a mathematical fallacy

OpenStudy (ghazi):

but i can prove it

OpenStudy (experimentx):

don't post answer ... i just need explanation http://projecteuler.net/problem=33

OpenStudy (ghazi):

 don't get irked please :)

OpenStudy (experimentx):

Woops!!

OpenStudy (ghazi):

lol

OpenStudy (experimentx):

funny!!

OpenStudy (experimentx):

i guess i understand what i means!!

OpenStudy (ghazi):

@experimentX can you please elaborate last line of your question?

OpenStudy (experimentx):

this isn't my question if it were mine i would have already solved it.

OpenStudy (experimentx):

i didn't expect the nine would be cancelled to give right answer.

OpenStudy (ghazi):

i know :)

OpenStudy (ghazi):

but in the example 30/50 cancel out zeroes you're gonna get that you're lookin for..sorry, let me try this

OpenStudy (experimentx):

9*9*9 possibilities ... if you get out trivial case ... 9*9*9 - 10 possible cases.

OpenStudy (ghazi):

no, i think there are exactly 4 possibilities and if they are expressed as lowest fraction you'll have denominator of it will be equal to ?? the other example is 16/64

OpenStudy (experimentx):

told ya .. don't post answer. this is supposed to be challenge Q.

OpenStudy (experimentx):

programming challenge ..

OpenStudy (ghazi):

well the answer is equal to = 2*5*4*2 and another example is 40/80

OpenStudy (ghazi):

\[\frac{ 30 }{ 50 },\frac{ 48 }{ 98 }, \frac{ 16 }{ 64 }, \frac{ 40 }{ 80 }\]

OpenStudy (experimentx):

zeros are trivial cases.

OpenStudy (ghazi):

sorry forgot that

OpenStudy (experimentx):

Yep .. got'em all

OpenStudy (ghazi):

??

OpenStudy (ghazi):

@experimentX what are those? i just got two

OpenStudy (anonymous):

there are 4 equation to solve\[\frac{ab}{cb}=\frac{a}{c}\]\[\frac{ba}{cb}=\frac{a}{c}\]\[\frac{ba}{bc}=\frac{a}{c}\]\[\frac{ab}{bc}=\frac{a}{c}\]if u wanna do it by hand of course :) these are easy to solve i think

OpenStudy (experimentx):

yep .. thanks for idea. looks like i hit jackpot. well ... the answers are 16 64 19 95 26 65 49 98

OpenStudy (ghazi):

no

OpenStudy (ghazi):

@experimentX 26/65 seems in correct

OpenStudy (experimentx):

they are multiples of 13

OpenStudy (ghazi):

exactly ..i was in haste ...nice you've got the jack pot ..cool